MAX-PLANCK-INSTITUT FÜR PHYSIK KOMPLEXER SYSTEME DRESDEN

International Workshop and Seminar on

Disordered Dynamical Systems Dresden

Workshop • 2 - 6 March, 1998 Seminar • 9 March - 9 April, 1998

Scientific Coordinator:

Günter Radons

University Kiel and FhG-IPA, Stuttgart

Max Planck Institute for Physics of Complex Systems

Nöthnitzer Strasse 38

D-01187 Dresden

Tel.: +49-351-871-2105; Fax: +49-351-871-2199 http://www.mpipks-dresden.mpg.de

WORKSHOP PROGRAM

Sunday, 1 March

Arrival

19:00 Welcome (Cafeteria at the Max Planck Institute)

Monday, 2 March

09:30 - 10:00 Peter Fulde (Dresden)

Günter Radons (Kiel)

Opening address

10:00 - 11:00 Leonid Bunimovich (Atlanta)

Space-time chaos in nonlinear wave equations

11:00 - 11:30 Coffee
11:30 - 12:30 Rudolf Hilfer (Stuttgart)

Correlated disorder and transport in porous media

12:30 - 14:00 Lunch
14:00 - 15:00 Henk van Beijeren (Utrecht)

Chaos and transport in low-density systems

15:00 - 16:00 Ramses van Zon (Utrecht)

Connection between dynamical systems theory and non-equilibrium statistical mechanics

16:00 - 16:30 Coffee
16:30 - 17:30 Harald Posch (Vienna)

Fractal phase space structures in non-equilibrium systems

18:00  Supper

Tuesday, 3 March

09:00 - 10:00 Pierre Gaspard (Bruxelles )

Chaos and transport in diffusion-reaction systems

10:00 - 10:30 Coffee
10:30 - 11:30 Andreas Engel (Magdeburg)

Multifractral phase space of neural networks

11:30 - 12:30 Rainer Klages (Bruxelles)

Deterministic transport, fractal transport coefficients,

and random perturbations

12:30 - 14:00 Lunch
14:00 - 15:00 Klaus W. Kehr (Jülich)

Random walks in systems with uncorrelated disorder

15:00 - 16:00 Matthias Gundlach (Bremen)

Products of random rectangular matrices for random walks in random environments

16:00 - 16:30 Coffee
16:30 - 17:30 Peter Hänggi (Augsburg)

Reaction rates when barriers fluctuate

18:00  Supper

Wednesday, 4 March

09:00 - 10:00 Dieter Weiss (Regensburg)

Transport experiments in periodic and random electric and magnetic fields

10:00 - 10:30 Coffee
10:30 - 11:30 Herbert Spohn (München)

Magnetotransport of the 2D Lorentz gas

11:30 - 12:30 Andreas Knauf (Leipzig)

Motion in periodic potentials

12:30 - 14:00 Lunch
14:00 - 15:00 Tyll Krüger (Leipzig)

Self-organized criticality and dynamical systems

15:00 - 16:00 Ga’bor Vattay (Budapest)

Periodic orbit theory of chaotic, noisy and quantum systems

16:00 - 16:30 Coffee
16:30 - 17:30 Otto E. Rössler (Tübingen)

Reversibility-based irreversibility and evolution

18:00  Supper

Thursday, 5 March

09:00 - 10:00 Charles M. Newman (New York)

Stochastic dynamics for (disordered) lattice spin systems

10:00 - 10:30 Coffee
10:30 - 11:30 Jean-Marc Luck (Gif-sur-Yvette)

Aging in simple systems

11:30 - 12:30 Lester Coyle (Baltimore)

to be announced

12:30 - 14:00 Lunch
14:00 - 15:00 Theo Geisel (Göttingen)

Anamalous wave packet spreading and correlation decay

15:00 - 16:00 Pal Rujan (Oldenburg)

Phase transitions in randomly driven spin systems

16:00 - 16:30 Coffee
16:30 - 17:30 Tamas Tel (Budapest)

Random maps in modelling chaotic flows

19:00  Workshop Dinner

Friday, 6 March

09:00 - 10:00 Conrad Perez-Vincente (Barcelona)

Recent contributions on the Kuramoto model

10:00 - 10:30 Coffee
10:30 - 11:30 Heinz Horner (Heidelberg)

to be announced

11:30 - 12:30 Sergej Flach (Dresden)

Discrete breathers - localization without disorder

12:30 - 14:00 Lunch
14:00 - 15:00 Walter Zimmermann (Jülich)

Heterogeneties in pattern formation

15:00 - 16:00 Joachim Krug (Essen)

Phase transitions in disordered exclusion models

16:00 - 16:30 Coffee
16:30 - 17:30 Itamar Procaccia (Rehovot)

Instability and noise in flame propagations:

dynamics in the complex plane

18:00  Supper

Saturday, 7 March - Free

Sunday, 8 March - Excursions

SEMINAR PROGRAM - next page SEMINAR PROGRAM

March 9 - April 9, 1998

Speakers tentatively scheduled for the 1st and 2nd Seminar week include

Hailemariam Ambaye (Jülich) Connection b/n Disspersive Transport and statistics of extrem events

Johannes Berg ( Magdeburg) Statistical mechanics of zero-sum games

Tseren Davaadorj (Köln) Synchranisation of time-delay coupled chaotic relaxational oscillators

Reza Ejtehadi (Tehran) Shape transition of polyampholytes in two dimensions

Andreas Engel (Magdeburg) to be announced

Hans Groeneveld (Utrecht) to be announced

George Guria (Moscow) Spatio-temporal dynamics of fibrin polymerization in the process of blood coagulation

Moma Jovanovic (Nis) Transition to complex spatio-temporal regimes in nonlinear wave interactions in plasmas

Klaus W. Kehr (Jülich) Random walks in random environments

Rainer Klages (Bruxelles) Simple deterministic dynamical systems with drift: Negative currents and suppression of diffusion

Andreas Knauf (Leipzig) The number-theoretical spin chain

Heiko Rieger (Jülich) Aging in disordered systems

Pal Rujan (Oldenburg) Disordered spin models for information processing

Jan C. Stiller (Kiel) Dynamics of nonlinear oscillators with random interactions

Balint Toth (Budapest) to be announced

Martin Weigt(Magdeburg) Replica theory of one-dimensional disordered systems Multifractal phase space of neural networks

During the last Seminar week (April 1 - 9)

Professor Imre Kondor (Budapest)

will give a series of lectures on

’The field theory of spin glasses and/ or static chaos in spin glasses’